OpenStudy (kainui):

How could I intuitively figure out that \[\frac{df}{dt}=\frac{\partial f}{\partial x} \frac{dx}{dt}+\frac{\partial f}{\partial y} \frac{dy}{dt}\] I feel like the linear independence of the variables is what makes this work so nicely but I'm not sure I could fully just figure this out on the fly if I wanted to.

OpenStudy (anonymous):

I can try and help you

OpenStudy (kainui):

Sure sounds good

OpenStudy (anonymous):

One moment while I generate an solution.

OpenStudy (kainui):

I'm trying to do a transformation from x and y to r and theta to see what the derivative would be to see if I get any kind of insight out of it.

OpenStudy (anonymous):

Give me 5-7 more minutes

OpenStudy (anonymous):

When T is dependent on x and y i.e t(x,y) then on the differentiation of t is done using chain rule

OpenStudy (anonymous):

take x=rcos theta and y=rsin theta

OpenStudy (anonymous):

Can you tell me exactly what you need me to find?

OpenStudy (kainui):

Well I just did it but I mean I'm really just blithely following the rule I've been given. It's not really giving me any insight into why this is the derivative of a multivariable function. I mean without checking, it could just as easily be something like \[\frac{df}{dt}=\frac{\partial ^2 f}{\partial x \partial y}* \frac{dx}{dt}* \frac{dy}{dt} \] You see what I mean?

OpenStudy (anonymous):

Oh yeah I see

OpenStudy (anonymous):

Do you want to try and solve it?

OpenStudy (kainui):

Yeah, I'm doing that right now actually lol.

OpenStudy (anonymous):

Okay :) I'm going to be trying as well.

OpenStudy (kainui):

I'm sort of playing around and seeing what this differential equation can do for me \[f_{xy}x'y'=f_x x'+f_y y'\] so the left side is my proposed definition of what it should be and the right side is what it is already. For clarity, f(x,y) with x(t) and y(t). So they're independent variables. No y(x) thing here.

OpenStudy (anonymous):

Ohhh, I see what you're doing. I'll show you mines in one moment

OpenStudy (kainui):

I think it is kinda a hard nonlinear differential equation to solve with no general solution so... Although interesting it doesn't really get me anywhere lol. I don't know if I really know what you're doing.

OpenStudy (anonymous):

\[\frac{ 2t }{ 2t }=\frac{ 2t }{ 2x } \frac{ dn }{ dt }+\frac{ 2t }{ 2y }\frac{ dy }{ dt }\] Here t is an function of x & y \[\frac{ 2t }{ 2x}\]is the partial derivation occurs only for x and rest of the variables are constant But \[\frac{ dt}{ dx } \] is the implicit derivative which occurs for both variables x & y with respect to x.

OpenStudy (anonymous):

There is much difference between partial and implicit derivation

OpenStudy (kainui):

I'm kinda lost with the 2's and the n. I'm assuming that n should be an x and the 2s are supposed to be partials.

OpenStudy (anonymous):

That's an example.

OpenStudy (kainui):

You've lost me.

OpenStudy (anonymous):

Oh no!!

OpenStudy (kainui):

I gotta go for now, I'll be back either later tonight or tomorrow! =)

OpenStudy (anonymous):

Alright :) Have a nice evening